OpenStudy (anonymous):

The equation of the tangent line to the curve y=x+√x at the point (16,20) is y = ?

OpenStudy (dumbcow):

slope of tangent line is dy/dx at given point dy/dx = 1+1/(2sqrtx) at x=16 dy/dx = 1+1/8 dy/dx = 9/8 slope is 9/8 next solve for y-intercept by plugging in given point to equation: y=mx+b